\begin{thebibliography}{99}
\bibitem{Fe} Fej\'{e}r L., Untersuchungen uber Fouriersche Reihen, Math.
Annalen, 58 (1904), 501--569.

\bibitem{FS}Fridli S., Schipp F., Strong summability and Sidon type inequalities. Acta Sci. Math. (Szeged) 60 (1995),
 no. 1-2, 277--289.

\bibitem{Ga} Gabisonia O. D., On strong summability points for Fourier
series, Mat. Zametki. 5, 14 (1973), 615--626.
\bibitem{Gar} Garnett J. B., Bounded analitic functions, 1981. translated in Russian.



\bibitem{GGT}G\'at G., Goginava U., Tkebuchava G., Convergence in measure of logarithmic means of quadratical
partial sums of double Walsh-Fourier series. J. Math. Anal. Appl. 323 (2006), no. 1, 535--549.

\bibitem{Gl}Glukhov V. A., Summation of multiple Fourier series in multiplicative systems.
(Russian) Mat. Zametki 39 (1986), no. 5, 665--673.

\bibitem{GG}Goginava U., Gogoladze L., Strong approximation by Marcinkiewicz means of two-dimensional
Walsh-Fourier series. Constr. Approx. 35 (2012), no. 1, 1--19.

\bibitem{Gogi}Goginava U., The weak type inequality for the maximal operator of the Marcinkiewicz-Fej\'er means
of the two-dimensional Walsh-Fourier series. J. Approx. Theory 154 (2008), no. 2, 161--180.


\bibitem{Gog} Gogoladze L. D., On strong summability almost everywhere.
(Russian) Mat. Sb. (N.S.) 135(177) (1988), no. 2, 158--168, 271; translation
in Math. USSR-Sb. 63 (1989), no. 1, 153--16


\bibitem{Gog2} Gogoladze L. D., Strong means of Marcinkiewicz type. (Russian) Soobshch. Akad. Nauk Gruzin.
SSR 102 (1981), no. 2, 293--295.


\bibitem{H-L} Hardy G. H., Littlewood J. E., Sur la series de Fourier
d'une fonction a carre sommable, Comptes Rendus (Paris) 156 (1913),
1307--1309.

\bibitem{Ka} Karagulyan G. A., Everywhere divergent $\Phi$-means of Fourier
series. (Russian) Mat. Zametki 80 (2006), no. 1, 50--59; translation in
Math. Notes 80 (2006), no. 1-2, 47--56

\bibitem{KrRu} Krasnoselski M. A., Rutitski Ya. B., Convex functions and Orlicz spaces, Moscow, 1958.(Russian)

\bibitem{Kon} Konyagin S. V., On the divergence of subsequences of partial
sums of multiple trigonometric Fourier series, Trudy MIAN 190 (1989),
102--116.

\bibitem{Le} Lebesgue H., Recherches sur la sommabilite forte des series de
Fourier, Math. Annalen 61 (1905), 251--280.

\bibitem{lei} Leindler L., Strong approximation by Fourier series, Akademiai Kiado, Budapest, 1985.

\bibitem{Ma} Marcinkiewicz J., Sur la sommabilit\'{e} forte de s\'{e}ries de
Fourier. (French) J. London Math. Soc. 14, (1939).162--168.

\bibitem{Ma2} Marcinkiewicz J., Sur une methode remarquable de sommation des
series doublefes de Fourier. \textit{Ann. Scuola Norm. Sup. Pisa}, 8(1939),
149--160.

\bibitem{Os} Oskolkov K. I., Strong summability of Fourier series. (Russian)
Studies in the theory of functions of several real variables and the
approximation of functions. Trudy Mat. Inst. Steklov. 172 (1985), 280--290.

\bibitem{Ro} Rodin, V. A., The space BMO and strong means of Fourier series.
Anal. Math. 16 (1990), no. 4, 291--302.

\bibitem{Sch} Schipp F., On the strong summability of Walsh series.
Dedicated to Professors Zolt\'{a}n Dar\'{o}czy and Imre K\'{a}tai. Publ.
Math. Debrecen 52 (1998), no. 3-4, 611--633.

\bibitem{S}Sj\"{o}lin P., Convergence almost everywhere of certain singular integrals and multiple Fourier
series. Ark. Mat. 9, 65--90. (1971).

\bibitem{tot} Totik V., On the strong approximation of Fourier series. Acta Math. Acad. Sci. Hungar.
35 (1980), no. 1-2, 151--172.

\bibitem{Zh} Zhizhiashvili L. V., Generalization of a theorem of Marcinkiwicz,
 \textit{Izvest.AN\ USSR, ser. matem. } 32(1968), 1112--1122 (Russian).

\bibitem{Zy} Zygmund A., On the convergence and summability of power series
on the circle of convergence II., Proc. London Math Soc. 47 (1941), 326--350.

\bibitem{Zy2} Zygmund A., Trigonometric series. Cambridge University Press,
Cambridge, 1959.

\bibitem{Wa} Wang, Kun Yang. Some estimates for the strong approximation of continuous periodic
functions of the two variables by their sums of Marcinkiewicz type. (Chinese) Beijing Shifan Daxue Xuebao 1981, no. 1, 7--22.

\bibitem{We} Weisz F., Strong Marcinkiewicz summability of multi-dimensional Fourier series. Ann. Univ. Sci.
Budapest. Sect. Comput. 29 (2008), 297--317.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage[active]{srcltx}
\usepackage{t1enc}
\usepackage[latin2]{inputenc}
\usepackage{verbatim}
\usepackage{amsmath,amsfonts,amssymb,amsthm}
\usepackage[mathcal]{eucal}
\usepackage{enumerate}
\usepackage[centertags]{amsmath}
\usepackage{graphics}
\usepackage[active]{srcltx}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\newtheorem*{OldTheorem}{Theorem A}
\newtheorem*{OldTheorem2}{Theorem B}
\newtheorem*{OldTheorem3}{Theorem C}
\newtheorem*{OldTheorem4}{Theorem D}

\def\bmo{{\rm BMO\,}}

\def\ZR{\ensuremath{\mathbb R}}

\def\ZT{\ensuremath{\mathbb T}}
\def\ZI{\ensuremath{\mathbb I}}
\def\ZB{\ensuremath{\mathcal B}}

\def\md#1#2\emd{
\ifx0#1\begin{equation*} #2 \end{equation*}\fi
\ifx1#1\begin{equation}#2\end{equation}\fi
\ifx2#1\begin{align*}#2\end{align*}\fi
\ifx3#1\begin{align}#2\end{align}\fi
\ifx4#1\begin{gather*}#2\end{gather*}\fi
\ifx5#1\begin{gather}#2\end{gather}\fi
\ifx6#1\begin{multline*}#2\end{multline*}\fi
\ifx7#1\begin{multline}#2\end{multline}\fi
\ifx8#1\begin{multline*}\begin{split}#2\end{split}\end{multline*}\fi
\ifx9#1\begin{multline}\begin{split}#2\end{split}\end{multline}\fi
}
\newcommand {\e }[1]{(\ref{#1})}

\newcommand {\trm }[1]{Theorem \ref{#1}}

\begin{document}

\author{U. Goginava, L. Gogoladze and G. Karagulyan}
\title[On the Exponential Almost Everywhere Summability]{BMO-estimation and Almost Everywhere Exponential Summability of
Quadratic Partial Sums of Double Fourier Series%
}
\address{U. Goginava, Department of Mathematics, Faculty of Exact and Natural
Sciences, Iv. Javakhishvili Tbilisi State University, Chavcha\-vadze str. 1, Tbilisi 0128,
Georgia}
\email{zazagoginava@gmail.com}

\address{L. Gogoladze, Department of Mathematics, Faculty of Exact and Natural
Sciences, Iv. Javakhishvili Tbilisi State University, Chavcha\-vadze str. 1, Tbilisi 0128,
Georgia}
\email{lgogoladze1@hotmail.com}

\address{G. Karagulyan, Institue of Mathematics of Armenian National Academy
of Science, Bughramian Ave. 24/5, 375019, Yerevan, Armenia}
\email{g.karagulyan@yahoo.com}

\maketitle

\begin{abstract}
It is proved a $\bmo$-estimation for quadratic partial sums of two-dimensional
Fourier series from which it is derived an almost everywhere exponential
summability of quadratic partial sums of double Fourier series.
\end{abstract}

\medskip

\footnotetext{%
2010 Mathematics Subject Classification: 40F05, 42B08
\par
Key words and phrases: Fourier series, Strong Summability, Quadratic sums.

\par
The research of U. Goginava was supported by Shota Rustaveli National
Science Foundation grant no.31/48 (Operators in some function spaces and their applications in
Fourier analysis)}

\section{Introduction}

Let $\mathbb{T}:=[-\pi ,\pi )=\ZR/2\pi$ and $\mathbb{R}:=\left( -\infty ,\infty
\right) $. We denote by $L_{1}\left( \mathbb{T}\right) $ the class of all
measurable functions $f$ on $\mathbb{R}$ that are $2\pi $-periodic and
satisfy
\begin{equation*}
\left\Vert f\right\Vert _{1}:=\int\limits_{\mathbb{T}}\left\vert
f\right\vert <\infty .
\end{equation*}%
The Fourier series of the function $f\in L_{1}\left( \mathbb{T}\right) $
with respect to the trigonometric system is the series%
\begin{equation}
\sum_{n=-\infty }^{\infty }\widehat{f}\left( n\right) e^{inx},
\label{fourier}
\end{equation}%
where
\begin{equation*}
\widehat{f}\left( n\right) :=\frac{1}{2\pi }\int\limits_{\mathbb{T}}f\left(
x\right) e^{-inx}dx
\end{equation*}%
are the Fourier coefficients of $f$.

Denote by $S_{n}(x,f)$ the partial sums of the Fourier series of $f$ and let
\begin{equation*}
\sigma _{n}(x,f)=\frac{1}{n+1}\sum_{k=0}^{n}S_{k}(x,f)
\end{equation*}%
be the $(C,1)$ means of (\ref{fourier}). Fej\'er \cite{Fe} proved that $\sigma
_{n}(f)$ converges to $f$ uniformly for any $2\pi $-periodic continuous
function. Lebesgue in \cite{Le} established almost everywhere convergence of
$(C,1)$ means if $f\in L_{1}(\mathbb{T})$. The strong summability problem,
i.e. the convergence of the strong means
\begin{equation}
\frac{1}{n+1}\sum\limits_{k=0}^{n}\left\vert S_{k}\left( x,f\right) -f\left(
x\right) \right\vert ^{p},\quad x\in \mathbb{T},\quad p>0,  \label{Hp}
\end{equation}%
was first considered by Hardy and Littlewood in \cite{H-L}. They showed that
for any $f\in L_{r}(\mathbb{T})~\left( 1<r<\infty \right) $ the strong means
tend to $0$ a.e., if $n\rightarrow \infty $. The trigonometric Fourier
series of $f\in L_{1}(\mathbb{T})$ is said to be $\left( H,p\right) $%
-summable at $x\in T$, if the values \e{Hp} converge to $0$ as $n\rightarrow
\infty $. The $\left( H,p\right) $-summability problem in $L_{1}(\mathbb{T})$
has been investigated by Marcinkiewicz \cite{Ma} for $p=2$, and later by
Zygmund \cite{Zy2} for the general case $1\leq p<\infty $. K.~I.~Oskolkov in
\cite{Os} proved the following

\begin{OldTheorem}
Let $f\in L_{1}(\mathbb{T})$ and let $\Phi $ be a continuous positive convex
function on $[0,+\infty )$ with $\Phi \left( 0\right) =0$ and
\begin{equation}  \label{a1}
\ln \Phi \left( t\right) =O\left( t/\ln \ln t\right) \text{ \ \ \ }\left(
t\rightarrow \infty \right) .
\end{equation}
Then for almost all $x$%
\begin{equation}
\lim\limits_{n\rightarrow \infty }\frac{1}{n+1}\sum\limits_{k=0}^{n}\Phi
\left( \left\vert S_{k}\left( x,f\right) -f\left( x\right) \right\vert
\right) =0.  \label{osk}
\end{equation}
\end{OldTheorem}

It was noted in \cite{Os} that V.~Totik announced the conjecture that (\ref%
{osk}) holds almost everywhere for any $f\in L_{1}(\mathbb{T})$, provided
\begin{equation}  \label{a2}
\ln \Phi \left( t\right) =O\left( t\right) \quad \left( t\rightarrow \infty
\right).
\end{equation}
In \cite{Ro} V.Rodin proved

\begin{OldTheorem2}
Let $f\in L_{1}(\mathbb{T})$. Then for any $A>0$
\begin{equation*}
\lim\limits_{n\rightarrow \infty }\frac{1}{n+1}\sum\limits_{k=0}^{n}\left(
\exp \left( A\left\vert S_{k}\left( x,f\right) -f\left( x\right) \right\vert
\right) -1\right) =0
\end{equation*}%
for a. e. $x\in \mathbb{T}$.
\end{OldTheorem2}

G.~Karagulyan \cite{Ka} proved that the following is true.

\begin{OldTheorem3}
Suppose that a continuous increasing function $\Phi :[0,\infty )\rightarrow
\lbrack 0,\infty ),\Phi \left( 0\right) =0$, satisfies the condition%
\begin{equation*}
\limsup_{t\rightarrow +\infty }\frac{\log \Phi \left( t\right) }{t}=\infty .
\end{equation*}%
Then there exists a function $f\in L_{1}(\mathbb{T})$ for which the relation%
\begin{equation*}
\limsup_{n\rightarrow \infty }\frac{1}{n+1}\sum\limits_{k=0}^{n}\Phi \left(
\left\vert S_{k}\left( x,f\right) \right\vert \right) =\infty
\end{equation*}%
holds everywhere on $\mathbb{T}$.
\end{OldTheorem3}

In fact, Rodin in \cite{Ro} has obtained a $\bmo$ estimate for the partial
sums of Fourier series and his theorem stated above is obtained from that
estimate by using John-Nirenberg theorem. Recall the definition of $\bmo[0,1]$
space. It is the Banach space of functions $f\in L_{1}[0,1]$ with the norm

\begin{equation*}
\Vert f\Vert _{\bmo}=\mathfrak{R}(f)+\left\vert \int_{0}^{1}f(t)dt\right\vert
\end{equation*}%
where
\begin{equation*}
\mathfrak{R}(f)=\sup_{I}(|f-f_{I}|)_{I}, f_{I}=\frac{1}{|I|}\int_{I}f(t)dt
\end{equation*}%
and the supremum is taken over all intervals $I\subset \lbrack 0,1]$ (\cite%
{Gar}, chap. 6). Let $\{\xi _{n}:\,n=0,1,2,\ldots \}$ be an arbitrary
sequence of numbers. Taking $\delta _{k}^{n}=[k/{(n+1)},(k+1)/{(n+1)}]$, we
define

\begin{equation*}
\bmo\left[ \xi _{n}\right] =\sup\limits_{0\leq n<\infty }\left\Vert
\sum\limits_{k=0}^{n}\xi _{k}\mathbb{I}_{\delta _{k}^{n}}\left( t\right)
\right\Vert _{\bmo}
\end{equation*}%
where $\mathbb{I}_{\delta _{k}^{n}}(t)$ is the characteristic function of $%
\delta _{k}^{n}$. Notice that the expressions
\begin{equation}
\bmo\left[ \widetilde{S}_{n}\left( x,f\right) \right] ,~\ \ \bmo\left[
S_{n}\left( x,f\right) \right] ,~\ f\in L_{1}\left( \mathbb{T}\right)
,x\in \mathbb{T}  \label{e}
\end{equation}%
define a sublinear operators, where $\widetilde{S}_{n}\left( x,f\right) $ is
the conjugate partial sum. The following theorem is proved by Rodin in \cite%
{Ro}.

\begin{OldTheorem4}
The operators (\ref{e}) are of weak type $(1,1)$, i.e. the inequalities
\begin{equation}\label{bmo1}
|\{x\in \mathbb{T}:\,\bmo\mathrm{\,}[S_{n}(x,f)]>\lambda \}|\leq \frac{c}{%
\lambda }\int_{\mathbb{T}}|f(t)|dt
\end{equation}
and
\begin{equation}\label{bmo2}
|\{x\in \mathbb{T}:\,\bmo\mathrm{\,}[\widetilde{S}_{n}(x,f)]>\lambda \}|\leq
\frac{c}{\lambda }\int_{\mathbb{T}}|f(t)|dt
\end{equation}%
hold for any $f\in L_{1}(\mathbb{T})$.
\end{OldTheorem4}
In this paper we study the question of exponential summability of
quadratic partial sums of double Fourier series. Let $f\in L_{1}(\mathbb{T}%
^{2})$, be a function
with Fourier series
\begin{equation}
\sum_{m,n=-\infty }^{\infty }\widehat{f}\left( m,n\right) e^{i(mx+ny)},
\label{DF}
\end{equation}%
where
\begin{equation*}
\widehat{f}\left( m,n\right) =\frac{1}{4\pi ^{2}}\iint\limits_{\mathbb{T}%
^{2}}f(x,y)e^{-i(mx+ny)}dxdy
\end{equation*}%
are the Fourier coefficients of the function $f$. The rectangular partial
sums of (\ref{DF}) are defined as follows:
\begin{equation*}
S_{MN}\left( x,y,f\right) =\sum_{m=-M}^{M}\sum_{n=-N}^{N}\widehat{f}\left(
m,n\right) e^{i(mx+ny)}.
\end{equation*}%
We denote by $L\log L\left( \mathbb{T}^{2}\right) $ the class of measurable
functions $f$, with
\begin{equation*}
\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f|<\infty ,
\end{equation*}%
where $\log ^{+}u:=\mathbb{I}_{(1,\infty )}\log u.$ For quadratic partial sums
of two-dimensional trigonometric Fourier series Marcinkiewicz \cite{Ma2} has
proved, that if $f\in L\log L\left( \mathbb{T}^{2}\right) $, then%
\begin{equation*}
\lim\limits_{n\rightarrow \infty }\frac{1}{n+1}\sum\limits_{k=0}^{n}\left(
S_{kk}\left( x,y,f\right) -f\left( x,y\right) \right) =0
\end{equation*}%
for a. e. $\left( x,y\right) \in \mathbb{T}^{2}$. L.~Zhizhiashvili \cite{Zh}
improved this result showing that class $L\log L\left( \mathbb{T}^{2}\right)
$ can be replaced by $L_{1}\left( \mathbb{T}^{2}\right) $.

From a result of S.~Konyagin \cite{Kon} it follows that for every $%
\varepsilon >0$ there exists a function $f\in L\log ^{1-\varepsilon }\left(
\mathbb{T}^{2}\right) $ such that
\begin{equation}
\lim\limits_{n\rightarrow \infty }\frac{1}{n+1}\sum\limits_{k=0}^{n}\left%
\vert S_{kk}\left( x,y,f\right) -f\left( x,y\right) \right\vert \neq 0\text{
\ \ for a. e. }\left( x,y\right) \in \mathbb{T}^{2}.  \label{str}
\end{equation}

The main result of the present paper is the following.

\begin{theorem}\label{BMO}
If $f\in L\log L\left( \mathbb{T}^{2}\right) $,
then
\begin{eqnarray}
&&\left\vert \{(x,y)\in \mathbb{T}^{2}:\bmo[S_{nn}(f,x,y)]>\lambda \}\right\vert \\
\newline
&\leq &\frac{c}{\lambda }\left( 1+\iint\limits_{\mathbb{T}^{2}}|f|\log
^{+}|f)|\right)  \notag
\end{eqnarray}
for any $\lambda >0$, where $c$ is an absolute positive constant.
\end{theorem}

The following theorem shows that the quadratic sums of two-dimensional
Fourier series of a function $f\in L\log L\left( \mathbb{T}^{2}\right) $ are
almost everywhere exponentially summable to the function $f$. It will be
obtained from the previous theorem by using John-Nirenberg theorem.

\begin{theorem}
\label{exp} Suppose that $f\in L\log L\left( \mathbb{T}^{2}\right) $. Then
for any $A>0$%
\begin{equation*}
\lim\limits_{m\rightarrow \infty }\frac{1}{m+1}\sum\limits_{n=0}^{m}\left(
\exp \left( A\left\vert S_{nn}\left( x,y,f\right) -f\left( x,y\right)
\right\vert \right) -1\right) =0
\end{equation*}%
for a. e. $\left( x,y\right) \in \mathbb{T}^{2}$.
\end{theorem}

According to a Lemma of L. D. Gogoladze \cite{Gog}, this theorem can be
formulated in more general settings.

\begin{theorem}
Let $\psi:[0,\infty)\to [0,\infty)$ be a increasing function,  satisfying the conditions
\md4
\lim_{u\rightarrow 0 }\psi \left( u\right) =\psi \left( 0\right)=0,\newline
\limsup_{u\rightarrow \infty }\frac{\log \psi \left( u\right) }{u}<\infty .
\emd
Then for any $f\in L\log L\left( \mathbb{T} ^{2}\right)$ we have
\begin{equation*}
\lim\limits_{m\rightarrow \infty }\frac{1}{m+1}\sum\limits_{n=0}^{m}\psi
\left( \left\vert S_{nn}\left( x,y,f\right) -f\left( x,y\right) \right\vert
\right) =0
\end{equation*}
almost everywhere on $\mathbb{T} ^{2}$.
\end{theorem}

The results on Marcinkiewicz type strong summation for the Fourier series have
been investigated in \cite{FS,Ga,Gog2,Gl,GG,GGT,Gogi,lei,tot,Wa,We,Zh}

\section{Notations and lemmas}
The relation $a\lesssim b$ bellow stands for $a\leq c\cdot b$, where
$c$ is an absolute constant. The conjugate function of a given $f\in L_1(\ZT)$ is defined by
\md0
\tilde f(x)=\text{p.v.}\frac{1}{\pi }\int\limits_\ZT \frac{f(x+t)}{2\tan(t/2)}dt=\lim_{\varepsilon\to 0}\frac{1}{\pi }\int\limits_{\varepsilon<|t|<\pi} \frac{f(x+t)}{2\tan(t/2)}dt.
\emd
According to Kolmogorov's and Zygmund's inequalities (see \cite{Zy2}, chap. 7), we have%
\md3
|\{x\in \ZT:\, |\tilde f(x)|>\lambda\}|\lesssim \frac{\|f\|_{L_1(\ZT)}}{\lambda},\label{Kolmogorov}\\
\int\limits_{\ZT}|\tilde f(x)|dx\lesssim 1+\int\limits_{\mathbb{T}}|f(x)|\log ^{+}|f(x)|dx.
\label{Zygmund}
\emd
It will be used two simple properties of $\bmo$ norm below. First one says, if $\xi_n=c$, $n=1,2,\ldots $, then $\bmo\left[ \xi _{n}\right]=|c|$. The second one is, the bound
\md0
\bmo[\xi_n]\le 3\sup_n|\xi_n|.
\emd
We  shall consider the operators
\md0
U_n(x,f)=\text{p.v.}\frac{1}{\pi}\int\limits_\ZT\frac{\cos nt}{2\tan( t/2)}f(x+t)dt .
\emd
The following lemma is an immediate consequence of Theorem D.
\begin{lemma}
The inequality
\md0
|\{x\in \ZT:\, \bmo[U_n(x,f)]>\lambda\}| \lesssim \frac{\|f\|_{L_1(\ZT)}}{\lambda}
\emd
holds for any $f\in L_1(\ZT)$.
\end{lemma}
\begin{proof}
For the conjugate Dirichet kernel we have
\md9\label{Dir}
\tilde D_n(t)&=\frac{\cos(t/2)-\cos(n+1/2)t}{2\sin(t/2)}\\
&=\frac{1}{2\tan(t/2)}+\frac{\sin nt}{2}-\frac{\cos nt}{2\tan(t/2)}
\emd
and we get
\md8
\tilde S_n(x,f)&=\frac{1}{\pi}\int\limits_\ZT\tilde D_n(t)f(x+t)dt\\
&=\tilde f(x)+\frac{1}{2\pi}\int\limits_\ZT f(x+t)\sin ntdt-U_n(x,f).
\emd
Thus, applying simple properties of $\bmo$ norm, we obtain
\md0
\bmo[U_n(x,f)]\le |\tilde f(x)|+\frac{1}{2\pi}\int\limits_\ZT |f(t)|dt+\bmo\left[\tilde S_n(x,f)\right]
\emd
Applying the bound \e{Kolmogorov} and Theorem D, the last inequality completes the proof of lemma.
\end{proof}
We consider the square partial sums
\md1\label{a20}
S_{nn}\left( x,y,f\right) =\frac{1}{\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}%
\frac{\sin \left( n+1/2\right) t\sin \left( n+1/2\right) s}{4\sin \left(
t/2\right) \sin \left( s/2\right) }f\left( x+t,y+s\right) dtds
\emd
and their modification, defined by
\md0
S_{nn}^*\left( x,y,f\right) =\frac{1}{\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}%
\frac{\sin nt \sin ns}{4\tan \left(
t/2\right) \tan \left( s/2\right) }f\left( x+t,y+s\right) dtds.
\emd
\begin{lemma}
If $f\in L\log L(\ZT^2)$, then
\md0
\iint\limits_{\ZT^2}\sup_n\left|S_{nn}(x,y,f)-S_{nn}^*(x,y,f)\right|dxdy\lesssim  1+\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f|.
\emd
\end{lemma}
\begin{proof}
Substituting the expression for Dirichlet kernel
\md0
D_n(t)=\frac{\sin(n+1/2)t}{2\sin t/2}=\frac{\sin n t}{2\tan (t/2)} +\frac{\cos nt}{2}
\emd
in \e{a20}, we get
\md8
S_{nn}\left( x,y,f\right)&-S_{nn}^*(x,y,f)\\
&=\frac{1}{\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}
\frac{\sin n t\cdot\cos n s}{4\tan (t/2)}  f\left( x+t,y+s\right) dtds   \\
&+\frac{1}{\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}\frac{\cos n t\cdot\sin ns}{4\tan (s/2)} f\left( x+t,y+s\right) dtds  \\
&+\frac{1}{4\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}\cos n t\cdot\cos n s\cdot f\left( x+t,y+s\right) dtds  \\
&={S}_{nn}^{\left( 1\right) }\left(
x,y,f\right) +{S}_{nn}^{\left( 2\right) }\left( x,y,f\right) +{S}_{nn}%
^{\left( 3\right) }\left( x,y,f\right) .
\emd
It is clear, that
\begin{equation}\label{a17}
|S_{nn}^{(3)}(x,y,f)|\lesssim  \|f\|_{L^1(\ZT^2)} \lesssim
1+\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f| .
\end{equation}
Everywhere below the notation
\md0
\text{ p.v.}\iint_{\ZT^2}f(t,s)dtds
\emd
stands for either
\md0
\text{ p.v.}\int_{\ZT}\left(\text{p.v.}\int_\ZT f(t,s)dt\right)ds, \hbox { or } \text{ p.v.}\int_{\ZT}\left(\text{p.v.}\int_\ZT f(t,s)ds\right)dt
\emd
and in each cases we have equality of these two iterated integrals. To observe that we will need just the fact that $ f\in L\log L (\ZT)$ implies $\tilde f\in L_1(\ZT)$. Hence, making simple transformations and then changing the variables, we get
\md9\label{a14}
{S}_{nn}^{\left( 1\right) }&\left( x,y,f\right)  \\
&=\text{p.v.}\frac{1}{2\pi ^{2}}
\iint\limits_{\mathbb{T}^2}\frac{\sin n(t+s)}{2\tan (t/2)}
f\left( x+t,y+s\right) ds dt
\\
&+\text{p.v.}\frac{1}{2\pi ^{2}}
\iint\limits_{\mathbb{T}^2}\frac{\sin n (t-s)}{2\tan (t/2)}
 f\left( x+t,y+s\right) ds dt
\\
&=\text{p.v.}\frac{1}{2\pi ^{2}}
\iint\limits_{\mathbb{T}^2}\frac{\sin n u \cdot f\left( x+v,y+u-v\right)}{2\tan (v/2)}  dvdu\quad(u=t+s,\, v=t)
\\
&+\text{p.v.}\frac{1}{2\pi ^{2}}
\iint\limits_{\mathbb{T}^2}\frac{\sin n u \cdot f\left( x+v,y+v-u\right)}{2\tan (v/2)}  dvdu\quad(u=t-s,\, v=t)
\\
&=\frac{1}{2\pi }
\int\limits_{\mathbb{T}}\sin n u \left(\text{p.v.} \frac{1}{\pi }\int\limits_{\mathbb{T}} \frac{f\left( x+v,y+u-v\right)}{2\tan (v/2)}
dv\right)du
\\
&+\frac{1}{2\pi }
\int\limits_{\mathbb{T}}\sin n u \left(\text{p.v.} \frac{1}{\pi }\int\limits_{\mathbb{T}} \frac{f\left( x+v,y+v-u\right)}{2\tan (v/2)}
dv\right)du.
\emd
Observe, that the functions
\md8
&F_1(x,y,u)=\text{p.v.} \frac{1}{\pi }\int\limits_{\mathbb{T}} \frac{f\left( x+v,y+u-v\right)}{2\tan (v/2)}dv\\
&F_2(x,y,u)=\text{p.v.} \frac{1}{\pi }\int\limits_{\mathbb{T}} \frac{f\left( x+v,y+v-u\right)}{2\tan (v/2)}dv
\emd
are defined for almost all triples $(x,y,u)$. Moreover, we shall prove that
\md3\label{a15}
\iiint\limits_{\ZT^3}|F_i(x,y,u)|dxdydu\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f|, \quad i=1,2.
\emd
Consider the function $h(t,s,u):=f(t+s,t+u-s)$. Substituting $x=t+s$ and  $y=t-s$ in the expression of $F_1$, we get
\md0
F_1(t+s,t-s,u)=\text{p.v.} \frac{1}{\pi }\int\limits_{\mathbb{T}} \frac{h\left( t,s+v,u\right)}{2\tan (v/2)}dv.
\emd
Thus, first using the inequality \e{Zygmund} for variable $s$, then integrating by $t$ and $u$, we obtain
\md0
\iiint\limits_{\ZT^3}|F_1(t+s,t-s,u)|dsdtdu\lesssim 1+\iiint\limits_{\mathbb{T}^{3}}|h(t,s,u)||\log ^{+}|h(t,s,u)|dtdsdu.
\emd
After the change of variables $t=(x+y)/2$ and $s=(x-y)/2$ in the integrals, we get \e{a15} in the case $i=1$. The case $i=2$ may be proved similarly.
On the other hand, from \e{a14} it follows that
\md0
|S_{nn}^{(1)}\left( x,y,f\right)|\le \frac{1}{2\pi}\int\limits_\ZT|F_1(x,y,u)|du+\frac{1}{2\pi}\int\limits_\ZT|F_2(x,y,u)|du.
\emd
Combining this inequality with \e{a15}, we obtain
\md1\label{a16}
\iint\limits_{\ZT^2}\sup_n|S_{nn}^{(1)}\left( x,y,f\right)|dxdy\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f|
\emd
Similarly we can get the same bound for  $S_{nn}^{(2)}\left( x,y,f\right)$, which together with \e{a17} completes the proof of lemma.
\end{proof}

\section{Proof of Theorems}
\begin{proof}[Proof of Theorem \ref{BMO}]
From Lemma 2 we obtain
\md0
|S_{nn}(x,y,f)-S_{nn}^*(x,y,f)|\le \phi(x,y),\quad n=1,2,\ldots,
\emd
where the function $\phi(x,y)\ge 0$ satisfies the bound
\md0
\iint\limits_{\ZT^2}\phi(x,y)dxdy\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|f|\log ^{+}|f|.
\emd
Thus we get
\md0
\bmo[S_{nn}(x,y,f)]\le \bmo[S_{nn}^*(x,y,f)]+3\phi(x,y).
\emd
Hence, the theorem will be proved, if we obtain $\bmo$ weak $(1,1)$ estimate for modified partial sums.  We have
\md9\label{a1}
S_{nn}^*\left( x,y,f\right) &  \\
=&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}
^{2}}\frac{\cos n(t-s)\cdot f\left(
x+t,y+s\right)}{4\tan \left(
t/2\right) \tan \left( s/2\right) } dtds\notag \\
-&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}
^{2}}\frac{\cos n(t+s)\cdot f\left(
x+t,y+s\right)}{4\tan \left(
t/2\right) \tan \left( s/2\right) } dtds\notag \\
=&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}
^{2}}\frac{\cos nu \cdot f\left(
x+u+v,y+v\right)}{4\tan \left(
(u+v)/2\right) \tan \left( v/2\right) } dudv\quad (u=t-s,v=s)\notag \\
-&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}
^{2}}\frac{\cos nu \cdot f\left(
x+u+v,y-v\right)}{4\tan \left(
(u+v)/2\right) \tan \left( v/2\right) } dudv\quad (u=t+s,v=-s)\notag\\
=& I_n(x,y,f)-J_n(x,y,f).\notag
\emd
Using a simple and an important identity
\md7
\frac{1}{\tan ((u+v)/2)\tan \left( v/2\right) }=\\
\frac{1}{\tan(u/2)\tan(v/2)}-\frac{1}{\tan(u/2)\tan ((u+v)/2)}-1,
\emd
we obtain
\md8
I_n(x,y,f)\\
=&\text{p.v.}\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}^2}\frac{\cos nu\cdot f\left(x+u+v,y+v\right)}{4\tan(u/2)\tan(v/2) } dudv\\
-&\text{p.v.}\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}\frac{\cos nu\cdot f\left(x+u+v,y+v\right) }{4\tan(u/2)\tan((u+v)/2) }dudv\\
-&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}f\left(x+t,y+s\right) dtds\\
= &\text{p.v.}\frac{1}{2\pi }\int\limits_{\mathbb{T}}\frac{\cos nu}{2\tan(u/2) }\left(\text{p.v.} \frac{1}{\pi } \int\limits_{\mathbb{T}}\frac{f\left(x+u+v,y+v\right) } {2\tan(v/2 ) }dv\right)du\\
- &\text{p.v.}\frac{1}{2\pi }\int\limits_{\mathbb{T}}\frac{\cos nu}{2\tan(u/2) }\left(\text{p.v.}\frac{1}{\pi } \int\limits_{\mathbb{T}}\frac{f\left(x+u+v,y+v\right) } {2\tan((u+v)/2 ) }dv\right)du\\
-&\frac{1}{2\pi ^{2}}\iint\limits_{\mathbb{T}^{2}}f\left(t,s\right) dtds=I_n^{(1)}(x,y,f)-I_n^{(2)}(x,y,f)-I^{(0)},
\emd
where
\md1\label{a7}
|I^{(0)}|=\frac{1}{2\pi ^{2}}\left|\iint\limits_{\mathbb{T}^{2}}f\left(t,s\right) dtds\right|\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|f(x,y)|\log ^{+}|f(x,y)|dxdy.
\emd
Observe that
\md0
I_n^{(1)}(x,y,f)=\frac{1}{2}\cdot U_n(x,A(\cdot, y))
\emd
where
\md0
A(x,y)=\text{p.v.}\frac{1}{\pi } \int\limits_{\mathbb{T}}\frac{f\left(x+v,y+v\right) } {2\tan(v/2 ) }dv.
\emd
Denoting $g(t,s):=f(t+s,t-s)$ and substituting $x=t+s$ and  $y=t-s$ we get
\md0
A(t+s,t-s)=\text{p.v.}\frac{1}{\pi } \int\limits_{\mathbb{T}}\frac{g\left(t+v,s\right) } {2\tan(v/2 ) }dv.
\emd
Using the inequality \e{Zygmund} for variable $t$ and then integrating by $s$, we obtain
\md0
\iint\limits_{\ZT^2}|A(t+s,t-s)|dsdt\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|g(t,s)||\log ^{+}|g(t,s)|dtds.
\emd
After the changing back of variables $t=(x+y)/2$ and $s=(x-y)/2$ we get
\begin{equation}
\iint\limits_{\mathbb{T}^{2}}\left\vert A\left( x,y\right) \right\vert
dxdy\lesssim 1+\iint\limits_{\mathbb{T}^{2}}|f(x,y)|\log ^{+}|f(x,y)|dxdy.
\label{Zy1}
\end{equation}
Hence, applying the Lemma 1, we conclude
\md7\label{a4}
|\{(x,y)\in \mathbb{T}^2:\,\bmo[I_n^{(1)}(x,y,f)]>\lambda \}|\\
\lesssim \frac{1}{%
\lambda }\left(1+\iint\limits_{\mathbb{T}^{2}}|f(x,y)|\log ^{+}|f(x,y)|dxdy\right).
\emd
After the changing of variable $u+v\to \nu$ in the inner integral of the expression of $I_n^{(2)}(x,y,f)$ we get
\md0
I_n^{(2)}(x,y,f)=\text{p.v.}\frac{1}{2\pi }\int\limits_{\mathbb{T}}\frac{\cos nu}{2\tan(u/2) }\left(\text{p.v.}\frac{1}{\pi } \int\limits_{\mathbb{T}}\frac{f\left(x+\nu,y+\nu-u\right) } {2\tan(\nu/2 ) }d\nu\right)du,
\emd
and then analogously we can prove that
\md7\label{a5}
|\{(x,y)\in \mathbb{T}^2:\,\bmo[I_n^{(2)}(x,y,f)]>\lambda \}|\\
\lesssim \frac{1}{%
\lambda }\left(1+\iint\limits_{\mathbb{T}^{2}}|f(x,y)|\log ^{+}|f(x,y)|dxdy\right).
\emd
Hence, using \e{a7}, \e{a4} and \e{a5}, we obtain
\md6
|\{(x,y)\in \mathbb{T}^2:\,\bmo[I_n(x,y,f)]>\lambda \}|\\
\lesssim \frac{1}{%
\lambda }\left(1+\iint\limits_{\mathbb{T}^{2}}|f(x,y)|\log ^{+}|f(x,y)|dxdy\right).
\emd
Using the absolutely same process we may get the analogous estimate for $J_n(x,y,f)$ and therefore  for $S_{nn}^*(x,y,f)$.
The theorem is proved.
\end{proof}
Let $X$ be either $[0,1]$ or $\ZT^2$ and $L_{M}=L_{M}(X)$ is the Orlicz space of functions on $X$, generated by Young function $M$,
 i. e. $M$ is convex continuous even function such that $M\left( 0\right)
=0 $ and%
\begin{equation*}
\lim_{t\rightarrow 0+}\frac{M(t)}{t}=\lim_{t\rightarrow \infty }\frac{t}{M(t)%
}=0.
\end{equation*}%
It is well known that $L_{M}$ is a Banach space with respect to Luxemburg norm%
\begin{equation*}
\left\Vert f\right\Vert _{(M)}:=\inf \left\{ \lambda :\,\lambda
>0,\,\int\limits_X M\left( \frac{|f|}{\lambda }\right) \leq
1\right\} <\infty .
\end{equation*}
 We will need some basic
properties of Orlicz spaces (see \cite{KrRu} ).

1) According to a theorem from (\cite{KrRu}, chap. 2, theorem 9.5) we have%
\begin{equation}  \label{d1}
\left\Vert f\right\Vert _{(M)}\leq 1\Rightarrow \int\limits_X
M\left( |f|\right) \leq \left\Vert f\right\Vert _{(M)},
\end{equation}%

2) From this fact we may deduce, that%
\begin{equation}  \label{d3}
0,5\left( 1+\int\limits_XM\left( |f|\right) \right) \leq
\left\Vert f\right\Vert _{(M)}\leq  1+\int\limits_X
M\left( |f|\right)
\end{equation}%
provided $\left\Vert f\right\Vert _{(M)}=1$.

3) From the definition of norm $\|\cdot\|_{(M)}$ immediately follows that $|f(x)|\le |g(x)|$ implies $\|f\|_{(M)}\le \|g\|_{(M)}$. Besides, for any measurable set $E$ we have%
\begin{equation*}
\left\Vert \mathbb{I}_{E}\right\Vert _{(M)}=o\left( 1\right) \text{ as }%
\left\vert E\right\vert \rightarrow 0\text{ \ (\cite{KrRu}, (9.23)).}
\end{equation*}

4) If $M$ satisfies $\Delta _{2}$-condition, that is%
\begin{equation*}
M\left( 2t\right) \le cM\left( t\right) ,t>t_{0},
\end{equation*}%
and $X=\ZT^2$, then the set of two variable trigonometric polynomials on $\ZT^2$ is dense in $L_{M}$ (\cite{KrRu}, \S 10).

5) From (\ref{d1}) it follows that for any sequence of functions $f_{n}$ the
condition $\left\Vert f_{n}\right\Vert _{(M)}\rightarrow 0$ implies $%
\int\limits_XM\left( |f_{n}|\right) \rightarrow 0.$

\begin{proof}[Proof of \trm{exp}]
We will deal with two $M$-functions
\md4
\Phi(t)=t\log^+t,\\
\Psi(t)=\exp t-1.
\emd
We consider two Orlicz spaces $L_\Phi= L_\Phi(\ZT^2)$ and  $L_\Psi= L_\Psi(0,1)$. Combining \e{d3} with \trm{BMO}, we may obtain
\md1\label{d4}
|\{ ( x,y) \in \mathbb{T}^{2}:\,\bmo [S_{nn}(x,y,f)] >\lambda \}
\lesssim \frac{\|f\|_{(\Phi)}}{\lambda}.
\emd
Indeed, at first we deduce the case when $\|f\|_{(\Phi)}= 1$, then, using a linearity principle, we get the inequality in the general case.

The inequality
\md1\label{d5}
\|f\|_{(\Psi)} \lesssim \|f\|_{\bmo}
\emd
proved in \cite{Ro}. It is an immediate consequence of the John-Nirenberg theorem.
Denote
\md1\label{d6}
\ZB f(x,y)=\sup_{0\le n<\infty}\left\|\sum_{k=0}^nS_{kk}(x,y,f)\ZI_{\delta_k^n}(t)\right\|_{{(\Psi)}}.
\emd
Notice, that by the definition we have
\md0
\bmo [S_{nn}(f,x,y)]=\sup_{0\le n<\infty}\left\|\sum_{k=0}^nS_{kk}(x,y,f)\ZI_{\delta_k^n}(t)\right\|_{\bmo}.
\emd
So, taking into account  \e{d4} and \e{d5} we obtain
\md1\label{d7}
|\{ ( x,y) \in \mathbb{T}^{2}:\,\ZB f(x,y) >\lambda \}
\lesssim \frac{\|f\|_{(\Phi)}}{\lambda}.
\emd
On the other hand we have
\md6
\frac{1}{n+1}\sum_{k=0}^{n}(
\exp A|S_{kk}( x,y,f) -f( x,y)| -1) \\
=\frac{1}{n+1}\sum_{k=0}^n\Psi(A|S_{kk}( x,y,f) -f( x,y)|)\\
=\int_0^1\Psi\left(A\sum_{k=0}^n|S_{kk}(x,y,f)-f(x,y)|\ZI_{\delta_k^n}(t)\right)dt.
\emd
Thus, according the property 5) of Orlicz spaces, to prove the theorem it is enough to prove that
\md1\label{d8}
\left\|\sum_{k=0}^n(S_{kk}(x,y,f)-f(x,y))\ZI_{\delta_k^n}(t)\right\|_{{(\Psi)}}\to 0,
\emd
almost everywhere on $\ZT^2$ as $n\to\infty$, for any $f\in L_\Phi$. It is easy to observe, that \e{d8} holds if $f$ is a real trigonometric polynomial in two variables. Indeed, if $P(x,y)$ is a polynomial of degree $m$, then we have
\md0
S_{kk}(x,y,P)-P(x,y)\equiv 0,\quad  k\ge m.
\emd
Therefore, if $n\ge m$, then we get
\md0
\left|\sum_{k=0}^n(S_{kk}(x,y,P)-P(x,y))\ZI_{\delta_k^n}(t)\right|\le C\cdot\ZI_{[0,m/(n+1)]}(t),
\emd
where $C$ is a constant, depending on $P$. Then, applying the property 3) of Orlicz spaces, we conclude that \e{d8} holds if $f=P$. To prove the general case, we consider the set
\md7\label{d9}
G_\lambda=\{(x,y)\in\ZT^2:\\ \limsup_{n\to\infty}\left\|\sum_{k=0}^n(S_{kk}(x,y,f)-f(x,y))\ZI_{\delta_k^n}(t)\right\|_{{(\Psi)}}>\lambda\}.
\emd
To complete the proof of theorem, it enough to prove that $|G_\lambda|=0$ if $\lambda>0$. It is easy to check that $\Phi(t)$ satisfies the $\Delta_2$-condition. Therefore, according the property 4), we may chose a polynomial $P(x,y)$ such that
$\|f-P\|_{(\Phi)}<\varepsilon$. Using the definition of $(\Phi)$-norm, we get
\md0
\int\limits_{\ZT^2}\Phi\left(\left|\frac{f-P}{\varepsilon}\right|\right)<1.
\emd
From Chebishev's inequality, one can easily deduce
\md0
|\{(x,y)\in\ZT^2:\,|f(x,y)-P(x,y)|>\lambda\}|\le \frac{1}{\Phi(\lambda/\varepsilon)},\quad \lambda>0.
\emd
Thus, using \e{d7} for any $\lambda>0$ we get
\md8
|G&_\lambda|=|\{(x,y)\in\ZT^2:\\
&\limsup_{n\to\infty}\left\|\sum_{k=0}^n(S_{kk}(x,y,f-P)-f(x,y)+P(x,y))\ZI_{\delta_k^n}(dt)\right\|_{{(\Psi)}}>\lambda\}|\\
&\le |\{ \ZB (f-P)(x,y)+c|f(x,y)-P(x,y)|>\lambda\}|\\
&\lesssim \frac{\|f-P\|_{(\Phi)}}{\lambda}+\frac{1}{\Phi(\lambda/\varepsilon)}\le \frac{\varepsilon}{\lambda}+\frac{1}{\Phi(\lambda/\varepsilon)}.
\emd
Since $\varepsilon>0$ may be taken sufficiently small, we conclude $|G_\lambda|=0$ if $\lambda>0$.
\end{proof}

{\bf Acknowledgement}. The authors would like to thank the referees for helpful suggestions.

\end{document}